Fermi National Laboratory
 

Operation Summary for July 29, 2011 - August 1, 2011

 

The Integrated Luminosity for the period from 7/25/11 to 8/1/11 was 17.71 pb-1 (inverse picobarns). During the same time period NuMI took 8.8E17 of beam.

Friday July 29

The day shift began with Operations recovering the Accelerator Complex from the power outage to bring the Main Substation (MSS) back into normal operational mode, with the Tevatron (TEV) off, with the Antiproton Source (Pbar) off, with the Recycler and Main Injector (MI) off, with Linac, and Booster off, with the MiniBooNE and Meson FTBF T-1008 experiments off, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) off.

During the day shift, operators reported that all CAMAC links had been restored. Operators prepared TEV F sector for access at 8:29 AM. Operators prepared CDF for access at 9:14 AM and Pbar for access at 9:30 AM. A NuMI person reported finding a problem with a NuMI Stub sump pump; FESS personnel investigated and adjusted the float switch. Booster RF technicians reported at 2:32 PM that they were swapping out the modulator for a Booster RF station (BRF14). Operators began ramping MI at 2:52 PM. Operations established beam to the Linac dump at 2:55 PM. Operators reported that the NuMI magnet E:V108 was ramping and that they had the NuMI beam permit. Operators reported that the MI beam permit was good. Operations established beam to the Booster dump at 3:46 PM.

During the evening shift, Operations established beam to the MI dump and then on to the NuMI target at 4:58 PM. Operators reported at 6:28 PM that they were unable to turn on the Switchyard power supplies because the TEV LCW system was down; a Water Works person estimated that the TEV LCW system might be up Sunday or Monday at the earliest. Operators successfully sent reverse protons through the Recycler at 6:57 PM. NuMI completed their target scans at 9:02 PM and began taking beam. At 11:35 PM, LRF1 tripped off; operators investigated, found a large LCW leak beneath the Power Amplifier, and called in Linac techs.

Saturday July 30

The midnight shift began with the TEV still in cryo system recovery mode, with Linac off due to LRF1 LCW leak repairs, and with no beam to anywhere.

During the midnight shift, a Linac tech completed his repairs on LRF1 at 2:03 AM; Linac resumed sending beam. At 3:09 AM, LRF1 tripped off for the first of three times in about 30 minutes. At 4:18 AM, the MIRF system tripped off due to a low LCW pressure indication; operators checked and could not find a problem. The RF system reset. At 5:10 AM, the MIRF system tripped off again; operators found that the nitrogen bottle for head pressure was empty. Operators replaced the bottle and beam returned. Operators prepared TEV F sector for access at 6:53 AM. Operators prepared Booster for access at 7:59 AM.

During the day shift, MTA personnel tried to began a study period at 8 AM, but controls problems halted the turn on. Water Works personnel reported at 8:40 AM that they had finished re-hosing MiniBooNE magnets in MI-12B. Operators turned on all the TEV pond water pumps at 10:11 AM per a Water Works request. Operations reestablished beam to the MI dump at 3:51 PM.

During the evening shift, Operations established beam to NuMI at 4:06 PM and to MiniBooNE at 4:13 PM. At 6:14 PM, Operations reported the first of eight kicker (I:KPS6N) trips in about 35 minutes; operators called in a Kicker person. Operators turned off the NuMI beam so the Kicker person could investigate the problem with KPS6N. Operators prepared the TEV F sector for access at 6:44 PM to investigate a vacuum valve problem. A Mechanical tech and operators investigated and found a compressed air leak.

Sunday July 32

The midnight shift began with no TEV beam due to cryo system recovery and beam to NuMI, MiniBooNE, and the Meson FTBF T-1008 experiments.

During the midnight shift, operators reported at 12:45 AM that the Mechanical tech had replaced a bad solenoid, but the valve still wouldn’t open. Another Mechanical tech came in to help. Operators reported at 3:03 AM that the F sector access and work was done and the valve opened. MiniBooNE resumed taking beam at 3:34 AM. At 3:56 AM, a MI Lambertson magnet (I:LAM52) began tripping off. Operators investigated and then called EE Support personnel to help. Also, problems with a MI power supply (I:H703) were keeping off the beam permit for the P1 Transfer line; operators called in an EE Support person. Operations established beam to NuMI at 3:58 AM.

During the day shift, operators reported at 8:51 AM problems with a MI quadrupole magnet power supply (I:QPS107) that was holding off beam; a fan on the power supply seemed to help. A CUB person reported at 9:20 AM that he was regenerating column #3. Operators reported at 10:23 AM that BRF16 tripped off for the first of six times in five minutes; a Booster person adjusted the station’s anode voltage. At 11:58 AM, operators reported that BRF16 went down again, operators bypassed the station, called in RF personnel, and prepared Booster for access. Operators reported at 1:39 PM that I:H703 was up and running.

During the evening shift, Booster RF personnel brought BRF16 back online at 4:46 PM and Booster resumed sending beam. Operations tried to establish beam to T-1008, but could not; operators called in Switchyard personnel. Switchyard personnel managed to resolve many problems and established beam to T-1008 at 11:31 PM.

Monday August 1

The midnight shift began with NuMI, MiniBooNE, and T-1008 taking beam.

During the midnight shift, at 5:09 AM, BRF16 tripped off for the first of four times in two hours; it reset. Operations disabled beam to T-1008 at 6:31 AM and prepared the TEV Transfer Hall for access.

The Plans for Monday

The plans for today are to continue with the TEV recovery, maintenance work, and sending beam to NuMI, MiniBooNE, and the Meson FTBF T-1008 experiments.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 08/01/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory